# Wet Hair: Global Symmetries in Entanglement Islands

### v1

@@ 5.2 The gravitational description
In this section, we describe the supergravity dual of the setup. presented before. It was constructed in DHoker:2007hhe ; DHoker:2007zhm (see also Aharony:2011yc ). We will see that the solution has two distinct regions: an $AdS_{5}\times{\bf{S}}^{5}$ region that is holographically dual to $\mathcal{N}=4$ SYM and an $AdS_{4}\times\mathbf{M}^{6}$ region (where $M_{6}$ is a six-dimensional manifold we specify below) that is the dual of the 3d $\mathcal{N}=4$ SCFT with the interpolating region should describe how the dual 3d and 4d systems couple in the holographic dual. This $AdS_{4}$ region caps off the $AdS_{5}$ and thus describes the Karch-Randall (KR) brane.
@@
The metric can be obtained explicitly, using the following intermediate expressions DHoker:2007hhe :
@@
We see that everything is constructed using the two harmonic functions $h_{1}$ and $h_{2}$ DHoker:2007hhe , which depend on the coordinate $w$ on the Riemann surface $\Sigma$. The functions $h_{1}$ and $h_{2}$ are expressed as:
@@
The functions Equ. (105), Equ. (106) parameterize the warping of the geometry along $w$ and can describe an $AdS_{5}$ with a KR ETW brane. In the limit of large $|w|$ the metric Equ. (100) describes $AdS_{5}\times S^{5}$ with $N$ units of RR 5-form flux, as can be checked using the explicit form of harmonic functions Equ. (105), Equ. (106) for $r\to\infty$ (see for example Aharony:2011yc ).
@@
We can see that the shape of this geometry is roughly the one of $AdS_{4}$ times a compact bag with some spikes at the positions of the 5-branes, which are given by $k_{a}$ and $l_{a}$. This bag has an $AdS_{5}\times S^{5}$ throat attached, which corresponds to the gravitational dual of the $D3$-branes ending on this web of 5-branes, see figure 5.3. This $AdS_{4}\times\mathbf{M}^{6}$ is the “fat” Karch-Randall brane in which the $AdS_{5}\times S^{5}$ ends. Notice that this string theory construction doesn’t have scale separation, so the length scales of the compact spaces are the same as their corresponding $AdS$ regions (see for example Bachas:2011xa ; Bachas:2018zmb ; Anastasi:2025puv . Moreover, this supergravity solution is not by itself UV complete due to the spikes Assel:2011xz . These spikes have a nice resolution in string theory as they should be replaced by the corresponding $5$-branes and the singularities just depict the fact that there are extra localized light degrees of freedom at these spikes that are captured by the low-energy worldvolume dynamics of the $5$-branes. Since the $D5$-branes and the NS5-branes are related to each other by $S$-duality, we choose to focus on the D3D3/$D5$ cases.
@@ 5.3 The Global Symmetries
In the model we considered in Sec. 3, we coupled a global symmetry in the bath to a global symmetry on its boundary through a marginal deformation Equ. (5), which is continuously tunable. However, in this string theory setup, as the solution is fully backreacted, the gravitational coupling between the ETW (the “bag” in figure ) and the bulk (the $AdS_{5}$ throat in figure ) is fixed by the number and positions of branes (see for instance Bachas:2018zmb ; Demulder:2022aij ; Huertas:2023syg ), it is not a marginal coupling, and we expect that the gauge symmetries in the bag and the throat will be coupled with the same strength as the gravity part. Therefore, weak coupling or weak leakage from the ETW part, will be given only by the scarcity of the bulk $CFT_{4}$ degrees of freedom, with respect those of the boundary $CFT_{3}$. This is a limit that we can obtain by increasing the number of 5-branes with respect to the semi-infinite D3-branes.
@@
The simplest way to obtain a gauge symmetry in the background Equ. (100) is to introduce additional spacetime filling branes in this background, for example a stack of M D7-branes, [^15] This stack of D7-branes fills the AdS4 and the Riemann surface $\Sigma$ with the other two directions wrapped on an $S^{2}$. as considered in Karch:2002sh (also used for the same purpose in Anastasi:2025puv ). If M is much smaller than N, we are in the probe limit, so we can take the approximation that the D7-branes don’t backreact on the background Equ. (100). Therefore, in the field theory description, we introduced a fundamental flavor group $U(M)$, with its global currents which are dual to the gauge fields living on the worldvolume of the spacetime filling D7-branes. This is a straightforward way to introduce a gauge field in the bulk of our string theory solution and is the realization in this setup of the toy model we considered in Sec. 4.
@@
In summary, we provided two types of realizations of the island setup global symmetry in the string theory setup we depicted in Sec. 5.1 and Sec. 5.2. This setup, as a potential string theory uplift of the Karch-Randall braneworld, has been used to provide explicit constructions of entanglement islands in string theory Uhlemann:2021nhu ; Demulder:2022aij ; Karch:2022rvr . These realizations of the global symmetry are aligned with our bottom-up studies in Sec. 3 and Sec. 4. Many interesting calculations regarding these realizations can be done on both the field theory side and the supergravity/string theory side, and we will come to this study in the near future.
6 Resolving the Harlow-Shagoulian Puzzle
With various concrete constructions of global symmetries in the island setup and a clear understanding of why global symmetry is consistent with entanglement islands and how it is encoded in the bath, we are ready to discuss the puzzle proposed by Harlow and Shagoulian in Harlow:2020bee . We first review the Harlow-Shagoulian puzzle and then provide the resolution of it based on our understanding of global symmetries in the island setup in Sec. 3 and Sec. 4.
6.1 The Harlow-Shagoulian Puzzle
Harlow and Shagoulian argued in Harlow:2020bee that a global symmetry is not consistent with entanglement islands. More specifically, they argued that there cannot exist operators inside the entanglement island that are charged under a global symmetry. Their argument is based on the following consideration. [^18] See Harlow:2018tng ; Harlow:2018jwu ; Bao:2024hwy for relevant considerations in different contexts. Let’s take an operator $\hat{O}_{q}^{\mathcal{I}}(x)$ with charge $q$ under a global symmetry inside the entanglement island $\mathcal{I}$ in a $d$-dimensional island setup. By the holographic duality, the global symmetry is supported in the bath and the CFTd-1 and by the holographic interpretation of entanglement islands, the global symmetry of the operator $\hat{O}_{q}^{\mathcal{I}}(x)$ should be encoded in the bath subregion $R$. Then due to the locality in the non-gravitational bath, one can write the symmetry transformation operator in the union of the bath and the CFTd-1 as
@@ 6.1 The Harlow-Shagoulian Puzzle
6.2 The Resolution of the Harlow-Shagoulian Puzzle
The resolution of the above puzzle is clear given our understanding that the global symmetry is always mixed with a bulk gauge symmetry and the fact that the AdS Goldstone boson is non-locally supported in both the AdS and in the bath. In the island setup, the operators inside the island are dressed using the Goldstone boson, so these dressed composite operators are gauge-invariant and have support also in the bath. Hence, the operator $\hat{O}_{q}^{\mathcal{I}}(x)$ localized inside the island from the perspective of the AdS necessarily intersects the entanglement wedge of a small bath subregion $R_{i}$. Thus, $U_{R_{i}}(g)$ can detect it. Note that the island operator $\hat{O}^{\mathcal{I}}_{q}(x)$ carries a lot of information and its bath dual is highly nonlocal and is not necessarily splittable as Equ. (111). We are suggesting that only the information of its global charge is splittable in the bath. Thus, the operator $\hat{O}^{\mathcal{I}}_{q}(x)$ doesn’t have to be inside the entanglement wedge of the small subregion $R_{i}$. Hence, there is no contradiction between such a global symmetry and the entanglement wedge reconstruction. This is readily seen in the Karch-Randall braneworld where it is clear that from the bulk perspective the global symmetry is in fact gauged and the Goldstone boson dressing is equivalent to a bulk smeared Wilson line dressing from the brane to the asymptotic boundary through the extra dimension. For the non-holographic island setups, one could think of the dressing to the bath as through a quantum extra dimension and the Goldstone boson is a microscopic wormhole operator Geng:2025rov .
@@ 7 Conclusions and Discussions
In this paper, we provided concrete evidence for the conjecture that the absence of global symmetry in quantum gravity is tied to unitarity and the realization of holography. Our strategy was to provide explicit constructions of global symmetries in controlled setups in which unitarity is violated. The examples we gave correspond to island setups that contain a gravitational asymptotically AdS spacetime coupled to a nongravitational bath on its asymptotic boundary. In such scenarios, in which the AdS contains an entanglement island, the AdS gravitational theory is not unitary. The holographic interpretation of entanglement islands states that information inside the islands is encoded in subregions in the nongravitational bath. Without an island, the who gravitational AdS region would be captured by its boundary, which however does not contain enough states to allow for a global symmetry in the AdS. On the other hand, black holes in island setups are no longer captured by the quantum system on its boundary but by the bath subregion that claims the island. The entanglement island overlaps the black hole interior. This resolves the apparent tension between a global symmetry and the Bekenstein-Hawking entropy formula. To put this conceptual understanding on firm ground, we provided explicit constructions of global symmetries in the island setup from both the bottom-up perspective and the top-down perspective. In these constructions, the previous understanding that gauge fields in the AdS are massive due to the bath induced spontaneous symmetry breaking in the island setup Geng:2020qvw ; Geng:2020fxl ; Geng:2021hlu ; Geng:2023zhq ; Geng:2023ynk ; Geng:2025rov played an essential role. General lessons from these constructions are that the global symmetry is always mixed with a spontaneously broken gauge symmetry and the charged operators in AdS under the global symmetry that are detectable by the bath can exist only inside the island and they are in fact nonlocally supported in both the AdS and the bath from the point of view of the full island setup. As a bonus, we realized a resolution of the Harlow-Shagoulian puzzle Harlow:2020bee from the above general lessons. Last but not least, we should notice that of entanglement islands, which is essential for our construction of global symmetry, universally exist in the island setup.
@@ Acknowledgments
We thank Safi Bahcall, Ning Bao, Liam A. Fitzpatrick, Tony Gherghetta, Jonathan Heckman, Luca Iliesiu, Yikun Jiang, Juan Maldacena, Yasunori Nomura, Matt Reece, John Terning, Huy Tran and Ángel Uranga for helpful discussions. HG, LR and DT are supported by the Gravity, Spacetime, and Particle Physics (GRASP) Initiative from Harvard University. JH is supported by a Simons Foundation’s fellowship through the Targeted Grant to Instituto Balseiro. AK is supported in part by DOE grant DE-SC0022021 and by a grant from the Simons Foundation (Grant 651678, AK).

### v2

@@ 5.2 The gravitational description
In this section, we describe the supergravity dual of the setup. presented before. It was constructed in D’Hoker et al. (2007b, a) (see also Aharony et al. (2011)). We will see that the solution has two distinct regions: an $AdS_{5}\times{\bf{S}}^{5}$ region that is holographically dual to $\mathcal{N}=4$ SYM and an $AdS_{4}\times\mathbf{M}^{6}$ region (where $M_{6}$ is a six-dimensional manifold we specify below) that is the dual of the 3d $\mathcal{N}=4$ SCFT with the interpolating region should describe how the dual 3d and 4d systems couple in the holographic dual. This $AdS_{4}$ region caps off the $AdS_{5}$ and thus describes the Karch-Randall (KR) brane.
@@
The metric can be obtained explicitly, using the following intermediate expressions D’Hoker et al. (2007b):
@@
We see that everything is constructed using the two harmonic functions $h_{1}$ and $h_{2}$ D’Hoker et al. (2007b), which depend on the coordinate $w$ on the Riemann surface $\Sigma$. The functions $h_{1}$ and $h_{2}$ are expressed as:
@@
The functions Equ. (105), Equ. (106) parameterize the warping of the geometry along $w$ and can describe an $AdS_{5}$ with a KR ETW brane. In the limit of large $|w|$ the metric Equ. (100) describes $AdS_{5}\times S^{5}$ with $N$ units of RR 5-form flux, as can be checked using the explicit form of harmonic functions Equ. (105), Equ. (106) for $r\to\infty$ (see for example Aharony et al. (2011)).
@@
We can see that the shape of this geometry is roughly the one of $AdS_{4}$ times a compact bag with some spikes at the positions of the 5-branes, which are given by $k_{a}$ and $l_{a}$. This bag has an $AdS_{5}\times S^{5}$ throat attached, which corresponds to the gravitational dual of the $D3$-branes ending on this web of 5-branes, see figure 5.3. This $AdS_{4}\times\mathbf{M}^{6}$ is the “fat” Karch-Randall brane in which the $AdS_{5}\times S^{5}$ ends. Notice that this string theory construction doesn’t have scale separation, so the length scales of the compact spaces are the same as their corresponding $AdS$ regions (see for example Bachas and Estes (2011); Bachas and Lavdas (2018); Anastasi et al. (2025). Moreover, this supergravity solution is not by itself UV complete due to the spikes Assel et al. (2011). These spikes have a nice resolution in string theory as they should be replaced by the corresponding $5$-branes and the singularities just depict the fact that there are extra localized light degrees of freedom at these spikes that are captured by the low-energy worldvolume dynamics of the $5$-branes. Since the $D5$-branes and the NS5-branes are related to each other by $S$-duality, we choose to focus on the D3D3/$D5$ cases.
@@ 5.3 The Global Symmetries
In the model we considered in Sec. 3, we coupled a global symmetry in the bath to a global symmetry on its boundary through a marginal deformation Equ. (5), which is continuously tunable. However, in this string theory setup, as the solution is fully backreacted, the gravitational coupling between the ETW (the “bag” in figure ) and the bulk (the $AdS_{5}$ throat in figure ) is fixed by the number and positions of branes (see for instance Bachas and Lavdas (2018); Demulder et al. (2023); Huertas and Uranga (2023)), it is not a marginal coupling, and we expect that the gauge symmetries in the bag and the throat will be coupled with the same strength as the gravity part. Therefore, weak coupling or weak leakage from the ETW part, will be given only by the scarcity of the bulk $CFT_{4}$ degrees of freedom, with respect those of the boundary $CFT_{3}$. This is a limit that we can obtain by increasing the number of 5-branes with respect to the semi-infinite D3-branes.
@@
The simplest way to obtain a gauge symmetry in the background Equ. (100) is to introduce additional spacetime filling branes in this background, for example a stack of M D7-branes, [^15] This stack of D7-branes fills the AdS4 and the Riemann surface $\Sigma$ with the other two directions wrapped on an $S^{2}$. as considered in Karch and Katz (2002) (also used for the same purpose in Anastasi et al. (2025)). If M is much smaller than N, we are in the probe limit, so we can take the approximation that the D7-branes don’t backreact on the background Equ. (100). Therefore, in the field theory description, we introduced a fundamental flavor group $U(M)$, with its global currents which are dual to the gauge fields living on the worldvolume of the spacetime filling D7-branes. This is a straightforward way to introduce a gauge field in the bulk of our string theory solution and is the realization in this setup of the toy model we considered in Sec. 4.
@@
In summary, we provided two types of realizations of the island setup global symmetry in the string theory setup we depicted in Sec. 5.1 and Sec. 5.2. This setup, as a potential string theory uplift of the Karch-Randall braneworld, has been used to provide explicit constructions of entanglement islands in string theory Uhlemann (2021); Demulder et al. (2023); Karch et al. (2022). These realizations of the global symmetry are aligned with our bottom-up studies in Sec. 3 and Sec. 4. Many interesting calculations regarding these realizations can be done on both the field theory side and the supergravity/string theory side, and we will come to this study in the near future.
6 Resolving the Harlow-Shaghoulian Puzzle
With various concrete constructions of global symmetries in the island setup and a clear understanding of why global symmetry is consistent with entanglement islands and how it is encoded in the bath, we are ready to discuss the puzzle proposed by Harlow and Shaghoulian in Harlow and Shaghoulian (2021). We first review the Harlow-Shaghoulian puzzle and then provide the resolution of it based on our understanding of global symmetries in the island setup in Sec. 3 and Sec. 4.
6.1 The Harlow-Shaghoulian Puzzle
Harlow and Shaghoulian argued in Harlow and Shaghoulian (2021) that a global symmetry is not consistent with entanglement islands. More specifically, they argued that there cannot exist operators inside the entanglement island that are charged under a global symmetry. Their argument is based on the following consideration. [^18] See Harlow and Ooguri (2018, 2019); Bao et al. (2025) for relevant considerations in different contexts and Heckman et al. (2024) for some comments on these considerations. Let’s take an operator $\hat{O}_{q}^{\mathcal{I}}(x)$ with charge $q$ under a global symmetry inside the entanglement island $\mathcal{I}$ in a $d$-dimensional island setup. By the holographic duality, the global symmetry is supported in the bath and the CFTd-1 and by the holographic interpretation of entanglement islands, the global symmetry of the operator $\hat{O}_{q}^{\mathcal{I}}(x)$ should be encoded in the bath subregion $R$. Then due to the locality in the non-gravitational bath, one can write the symmetry transformation operator in the union of the bath and the CFTd-1 as
@@ 6.1 The Harlow-Shaghoulian Puzzle
6.2 The Resolution of the Harlow-Shaghoulian Puzzle
The resolution of the above puzzle is clear given our understanding that the global symmetry is always mixed with a bulk gauge symmetry and the fact that the AdS Goldstone boson is non-locally supported in both the AdS and in the bath. In the island setup, the operators inside the island are dressed using the Goldstone boson, so these dressed composite operators are gauge-invariant and have support also in the bath. Hence, the operator $\hat{O}_{q}^{\mathcal{I}}(x)$ localized inside the island from the perspective of the AdS necessarily intersects the entanglement wedge of a small bath subregion $R_{i}$. Thus, $U_{R_{i}}(g)$ can detect it. Note that the island operator $\hat{O}^{\mathcal{I}}_{q}(x)$ carries a lot of information and its bath dual is highly nonlocal and is not necessarily splittable as Equ. (111). We are suggesting that only the information of its global charge is splittable in the bath. Thus, the operator $\hat{O}^{\mathcal{I}}_{q}(x)$ doesn’t have to be inside the entanglement wedge of the small subregion $R_{i}$. Hence, there is no contradiction between such a global symmetry and the entanglement wedge reconstruction. This is readily seen in the Karch-Randall braneworld where it is clear that from the bulk perspective the global symmetry is in fact gauged and the Goldstone boson dressing is equivalent to a bulk smeared Wilson line dressing from the brane to the asymptotic boundary through the extra dimension. For the non-holographic island setups, one could think of the dressing to the bath as through a quantum extra dimension and the Goldstone boson is a microscopic wormhole operator Geng (2025).
@@ 7 Conclusions and Discussions
In this paper, we provided concrete evidence for the conjecture that the absence of global symmetry in quantum gravity is tied to unitarity and the realization of holography. Our strategy was to provide explicit constructions of global symmetries in controlled setups in which unitarity is violated. The examples we gave correspond to island setups that contain a gravitational asymptotically AdS spacetime coupled to a nongravitational bath on its asymptotic boundary. In such scenarios, in which the AdS contains an entanglement island, the AdS gravitational theory is not unitary. The holographic interpretation of entanglement islands states that information inside the islands is encoded in subregions in the nongravitational bath. Without an island, the who gravitational AdS region would be captured by its boundary, which however does not contain enough states to allow for a global symmetry in the AdS. On the other hand, black holes in island setups are no longer captured by the quantum system on its boundary but by the bath subregion that claims the island. The entanglement island overlaps the black hole interior. This resolves the apparent tension between a global symmetry and the Bekenstein-Hawking entropy formula. To put this conceptual understanding on firm ground, we provided explicit constructions of global symmetries in the island setup from both the bottom-up perspective and the top-down perspective. In these constructions, the previous understanding that gauge fields in the AdS are massive due to the bath induced spontaneous symmetry breaking in the island setup Geng and Karch (2020); Geng et al. (2020, 2021); Geng (2023a, b, 2025) played an essential role. General lessons from these constructions are that the global symmetry is always mixed with a spontaneously broken gauge symmetry and the charged operators in AdS under the global symmetry that are detectable by the bath can exist only inside the island and they are in fact nonlocally supported in both the AdS and the bath from the point of view of the full island setup. As a bonus, we realized a resolution of the Harlow-Shaghoulian puzzle Harlow and Shaghoulian (2021) from the above general lessons. Last but not least, we should notice that of entanglement islands, which is essential for our construction of global symmetry, universally exist in the island setup.
@@ Acknowledgments
We thank Safi Bahcall, Ning Bao, Liam A. Fitzpatrick, Jonathan Heckman, Luca Iliesiu, Yikun Jiang, Juan Maldacena, Yasunori Nomura, Matt Reece, Edgar Shaghoulian, Huy Tran and Ángel Uranga for helpful discussions. HG, LR and DT are supported by the Gravity, Spacetime, and Particle Physics (GRASP) Initiative from Harvard University. JH is supported by a Simons Foundation’s fellowship through the Targeted Grant to Instituto Balseiro. AK is supported in part by DOE grant DE-SC0022021 and by a grant from the Simons Foundation (Grant 651678, AK).